# The NTU Partitioned Matching Game for International Kidney Exchange Programs

### v1

@@ Acknowledgement
The authors would like to thank Péter Biró and Zsuzsanna Jankó for the fruitful discussions on the topic. Gergely Csáji is supported by the Hungarian Scientific Research Fund, OTKA, Grant No. K143858 and by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation fund, financed under the KDP-2023 funding scheme (grant number C2258525). Tamás Király is supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – grant number ELTE TKP 2021-NKTA-62 and grant number K143858.

### v2

@@ 4.2 Strong core
5 Conclusions
We studied the NTU Partitioned Matching Game as a natural model for international kidney exchange programs (IKEPs), where players represent countries or institutions and utilities represent the number of patients matched. Our results show that in the special case where each player controls two vertices – the NTU Matching Game with Couples case – the weak core is always nonempty, and the existence of strong core solutions can be decided in polynomial time. Moreover, while optimizing over the weak core is NP-hard, we provide a polynomial-time algorithm to optimize over the strong core. We also prove that when the number of players is constant, all core-related decision and optimization problems are solvable in polynomial time. The latter findings may yield practically useful algorithms for IKEPs, where the number of participating countries is typically small, and they suggest core allocations as a viable and robust mechanism design goal in such settings.
Several questions remain open for future research. One direction is to explore different parameterized approaches for search and optimization in the weak or strong core (e.g. with graph-theoretic parameters), given its computational hardness. Another promising direction is to investigate dynamic or multi-round settings, reflecting how kidney exchanges are implemented in practice. Extending the model to incorporate more realistic constraints – such as donor compatibility scores, chain exchanges, or altruistic donors – could also bridge the gap between theory and application.
@@ Acknowledgement
The authors would like to thank Péter Biró and Zsuzsanna Jankó for the fruitful discussions on the topic, and the anonymous referees for their insightful comments. Gergely Csáji is supported by the Hungarian Scientific Research Fund, ADVANCED grant no. 150556, by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-2/2021 and by the Ministry of Culture and Innovation of Hungary from the National Research, Development and Innovation fund, financed under the KDP-2023 funding scheme (grant number C2258525). Tamás Király is supported by the Lendület Programme of the Hungarian Academy of Sciences – grant number LP2021-1/2021, by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund – ELTE TKP grant no. 2021-NKTA-62 and ADVANCED grant no. 150556.